Show that π‘ˆ(βˆ’π‘“, 𝑝) = βˆ’πΏ(𝑓, 𝑝) and 𝐿(βˆ’π‘“, 𝑝) = βˆ’π‘ˆ(𝑓, 𝑝)


1
Expert's answer
2021-10-12T13:40:17-0400

let p is partition on [a,b]

"p=\{a=x_0, x_1, x_2...x_{i-1}, x_i,...x_n=b \}"

is any partition on [a,b]


By definition lower and upper Riemman sum


"L(f,p)=\displaystyle\sum_{i=1}^nm_i\Delta x_i"


"U(f,p)=\displaystyle\sum_{i=1}^nm_i \Delta x_i"


where "m_i=int \{f(x)x\in[x_{i-1},x_i]\}"


"m_i=sup\{f(x):x\in[x_{i-1},x_i]\}"


"intrimum(-f(x))=-supremum \space f(x)"


"m_i=int \{-f(x)x\in[x_{i-1},x_i]\}"

"=-sup\{f(x):x\in[x_{i-1},x_i]\}"

"=-m_i"


"L(-f,p)=\displaystyle\sum_{i=1}^nm_i\Delta x_i"


"=-\displaystyle\sum_{i=1}^nm_i\Delta x_i"


"\ud835\udc3f(\u2212\ud835\udc53, \ud835\udc5d) = \u2212\ud835\udc48(\ud835\udc53, \ud835\udc5d).......proved"


now,

"m_i=sup\{-f(x):x\in[x_{i-1},x_i]\}"


"=-int \{f(x)x\in[x_{i-1},x_i]\}\\=-m_i"


"U(-f,p)=\displaystyle\sum_{i=1}^nm_i\Delta x_i"


"=-\displaystyle\sum_{i=1}^nm_i\Delta x_i"


"\ud835\udc48(\u2212\ud835\udc53, \ud835\udc5d) = \u2212\ud835\udc3f(\ud835\udc53, \ud835\udc5d)..........proved"